Circular No. 8997

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 U1 (McMILLAN) Robert S. McMillan, Lunar and Planetary Laboratory, University of

Arizona, reports his discovery of a comet with a 15′′ tail (spanning 20^{o}) toward the west-southwest and a coma of diameter ∼ 5′′ on CCD images taken with the Spacewatch 1.8-m f/2.7 reflector at Kitt Peak (discovery observation tabulated below). Four co-added 25-s unfiltered CCD images taken by A. R. Gibbs at the Mt. Lemmon 1.5-m reflector on Oct. 20.4 UT show a condensed coma of size 6′′ × 8′′, elongated in p.a. 250^{o}, with an additional “fuzziness” of diameter 10′′.

2008 UT  α_{2000}  δ_{2000}  Mag.
Oct. 19.19689  2^{h}03^{m}22.^{s}46  +11^{o}44′54′′.6  18.9

The available astrometry (including prediscovery observations with the Spacewatch 0.9-m f/3 reflector back to Sept. 20), the following elliptical orbital elements, and an ephemeris appear on MPEC 2008-U29.

T = 2008 May 12.8317 TT  ω = 310.^{o}3471
e = 0.375714  Ω = 36.4170 }2000.0
q = 2.523942 AU  i = 4.4150
a = 4.042929 AU n^{o} = 0.1212439 P = 8.13 years

V1309 SCORPII R. J. Rudy, D. K. Lynch, and R. W. Russell, Aerospace Corporation;

M. Sitko, Space Science Institute; C. E. Woodward, University of Minnesota; and C. Aspin, University of Hawaii, report that 0.8–5.5-μm spectroscopy and K-band imagery of the nova V1309 Sco (cf. IAUC 8972, 8976) on Oct. 4 and 16 UT using the Infrared Telescope Facility (+ Spex) shows very narrow emission lines of H I, He I, and Ca II sitting atop a strong continuum that closely matches that of a late M giant star. The spectrum includes strong molecular absorptions of CO, H_{2}O, and weaker features of TiO and VO, while the overall continuum shape is that of a cool star. Magnitudes derived from the spectroscopic observations are J = 6.7, H = 5.7, and K = 5.2. All of the observations would be consistent with V1309 Sco being a symbiotic nova (a nova consisting of a white dwarf and a late giant) except that the 2MASS images show no object close to the location of V1309 Sco with K < 11. If V1309 Sco has shed a massive envelope, such as V838 Mon or V4332 Sgr, its appearance at this epoch is very different. Additional observations of this interesting object are strongly encouraged.

2008 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green